OpenStudy (anonymous):

Determine the solution set of (3x + 1)2 - 100 = 0.

OpenStudy (anonymous):

its 3x +1 squared

OpenStudy (anonymous):

First decompose your (3x +1)^2. it will yield: \[(3x + 1)^2 - 100 = 0 \rightarrow 9x^2 + 6x + 1 - 100 = 0 \rightarrow 9x^2 + 6x - 99 = 0\] Use the quadratic formula. \[x = { -6 \pm \sqrt{36 + 4*9*99} \over 2*9 } = {-6 \pm \sqrt{3600} \over 18 } = -{1 \over 2} \pm {60 \over 18} = - {3 \over 6} \pm {8 \over 6}\] So your solutions are x = -11/6 and x = 5/6

OpenStudy (anonymous):

\[(3x+1)^{2}-10^{2} =0\] 3x+1+10 =0 or 3x+1-10=0 3x+11=0 or 3x-9=0 3x=-11 or 3x=9 x=-11/3 or x=3

OpenStudy (anonymous):

@soati \[(-6\pm60)/18 =(-1\pm10)/3=-11/3 or 9/3 = -11/3 or 3\]

OpenStudy (anonymous):

6/18 should be 1/3 rather 1/2 as in your calculations. Apart from this you did it good

OpenStudy (anonymous):

Youre right, my bad. I was focused on writting all the { x over y }'s hehe. Still my way works for a wider number of problems